# Sum frequency generation from real-time simulations in two-dimensional crystals

### v1

@@ 2.3 Field-Induced Second-Harmonic Generation
$P(2\omega^{\pm})=\hat{\chi}^{(3)}(2\omega\pm\nu;\nu,\omega_{,}\omega)E_{1}(\nu)E^{2}_{2}(\omega).$ (7)
If $E_{1}$ is static, i.e. $\nu=0$, the third order susceptibility extracted from this contribution to the polarization gives the FI-SHG. Similarly, if $\nu\ll\omega$, i.e $E_{1}$ is in the THz range, we have $2\omega\pm\nu\approx 2\omega$ and the expression in Eq. (7) can be rewritten as,
$P(2\omega^{\pm})\approx\hat{\chi}^{(3)}(2\omega;\nu,\omega,\omega)E_{1}(\nu)E^{2}_{2}(\omega).$ (8)
Then, summing the $\chi^{(3)}$ extracted by $P(2\omega^{+})$ and $P(2\omega^{-})$, one obtains the corresponding FI-SHG for low-frequency time-dependent pump fields.
@@ 3 Signal processing: nonlinear susceptibilities
Adding an extra field to the formalism presented in Ref. [11] to obtain Eq. (1) is straightforward. The challenging part, and the main contribution of this work, is finding feasible, accurate strategies for extracting the relevant nonlinear susceptibilities from the resulting polarization $\textbf{P}(t)$. A strategy based on the discrete-Fourier transform (DFT) is used in Ref. [11] (Sec. 3.1) for the case of one external monochromatic external field. This analysis can be extended to more external monochromatic fields (Sec. 3.2), but as the common period for two, or more, (commensurate) frequencies can be of several hundreds of fs, this implies very long and computationally costly simulations. We thus propose an alternative strategy based on the least square fit (LSF), which turns out to be as accurate as the DFT without requiring too long simulations.
@@ 3.1 One external monochromatic field
$\textbf{P}(t)=\sum_{n=-\infty}^{\infty}\bm{\mathrm{p}}^{(n)}\exp[-i\omega^{(n)}t]\mathrm{,}$ (9)
with $\omega^{(n)}=n\omega_{L}$. The complex Fourier components $\bm{\mathrm{p}}_{n}$ can be determined by truncating the Fourier series to an order $S$ and sampling $12S+$ 1 values $\textbf{P}_{i}\equiv\textbf{P}(t_{i})$ within a period $T_{L}=2\pi/\omega_{L}$. For each direction $\alpha$, this yields to
$\sum_{n=-S}^{S}\mathcal{F}^{(n)}_{i}p^{(n)}_{\alpha}=P_{\alpha,i}\quad i=1,2S+1\mathrm{,}$ (10)
@@
Figure 2: The time-dependent polarization (purple solid line) of h-BN calculated at the independent particle level with two electric fields ($\omega_{1}=0.2~{}\mathrm{eV}$, $\omega_{2}=3~{}\mathrm{eV}$). The signal can be divided into two regions: an initial “equilibration” region (up to $t\gg\gamma_{\mathrm{deph}}$, here $t=50~{}\mathrm{fs}$) during which the system’s eigenfrequencies are suppressed by dephasing and a region where Eq. (12) holds. In the second region the polarization is logarithmic sampled (teal dots) within a converged time window of 15 fs, smaller than the fundamental period of 20.678 fs of the signal. This sampling time is sufficient to correctly determine the Fourier coefficients by the nonlinear least square fit (LSF) as verified by reconstructing the polarization (lavender dashed line) within the fundamental period using the Fourier coefficients and the truncated Eq. (12).
@@ 3.2 Two external monochromatic fields
The resulting polarization $\textbf{P}(t)$ is periodic with period $T=2\pi/\omega_{0}$ and hence can be written as a Fourier series as in Eq. (9), in terms of the harmonics of $\omega_{0}$. As the frequencies of the external fields, are multiple of the fundamental frequency, $\omega_{1}=M\omega_{0}$ and $\omega_{2}=N\omega_{0}$, then the $|M\pm N|$ harmonics correspond to the SFG/DFG. When it comes to setting up the system of linear equations Eq. (10), the sum over the harmonics must be truncated to an appropriate $S$ to include these processes (that is $S\gtrapprox M+N$). When compared with a single external field, the dimension of system of linear equations is significantly larger. More critically, $T$ can be very long for any pair of frequencies with more than one decimal (see Fig. 3). Depending on the ratio of $\omega_{1}$ and $\omega_{2}$, $T$ can be orders of magnitude larger than the typical laser periods in the near-infrared to near-UV range ($\approx 1-5$ fs). As the computational expense increases linearly with simulation time, using these approaches to process the polarization signal from two monochromatic fields, may lead to computationally intensive simulations.
Alternatively, we can expand the polarization $\textbf{P}(t)$ as the product of two Fourier series, one in terms of the harmonics of $\omega_{1}$ and the other in terms of the harmonics of $\omega_{2}$,
$\textbf{P}(t)=\sum_{n,m=-\infty}^{\infty}\bm{\mathrm{C}}^{(n,m)}\mathrm{exp}\left[-i(n\omega_{1}+m\omega_{2})t\right]\mathrm{,}$ (12)
@@
$\sum_{n=-S_{1}}^{S_{1}}\sum_{m=-S_{2}}^{S_{2}}\mathcal{F}_{i}^{(n,m)}C^{(n,m)}_{\alpha}=P_{\alpha,i}\mathrm{,}$ (13)
where $\mathcal{F}_{i}^{(n,m)}\equiv\exp[-i(n\omega_{1}+m\omega_{2})t_{i}]$ and $S_{1},S_{2}$ are the maximum number of harmonics considered for each external field. Compared to the DFT with one external field, it is straightforward to identify the relevant coefficients, for example, $n=1,m=1$ gives the SFG and $n=2,m=\pm 1$ the FI-SHG (choosing $\omega_{2}=\Omega$). On the other hand, generally, the system of linear equations in Eq. (13) is ill-conditioned. To solve Eq. (13), we calculate the Moore-Penrose inverse [23, 24], also called pseudoinverse, of $\mathcal{F}_{i}^{(n,m)}$ by using the singular value decomposition (SVD). This approach (that we will refer to as SVD) allows to work with a much smaller sampling time than $T$, so to significantly reduce the time from the simulations when compared with the DFT.
A further approach to obtain the Fourier coefficients is by solving a nonlinear least squares problem [25]. This consists in finding the set of $C^{(n,m)}_{\alpha}$ that minimizes:
$\sum_{i=1}^{N}|P_{\alpha}(t_{i})-\bar{P}_{\alpha}(t_{i})|^{2},$ (14)
where $N$ is the number of sampling points and $\bar{\textbf{P}}$ is the Fourier series of Eq. (12) truncated to an order of $S$. Similarly to the SVD-based approach, the main advantage of the LSF over the DFT is that it is sufficient to sample part of the period $T$ to find the Fourier coefficient accurately. For instance, for h-BN, Fig. 2 compares the polarization from the simulation with that reconstructed from Eq. (12) with the coefficients obtained from LSF. Although only 15 fs are sampled, the LSF provides an accurate result.
The accuracy and performance of the three approaches are contrasted in Fig. 3 for the SFG of h-BN. Using the approaches based on the SVD and the LSF allows to cut simulation time by a factor 26 compared to the DFT.
Figure 3: SFG of h-BN with a pump frequency $\omega_{2}=3~{}\mathrm{eV}$ obtained at the IPA level using the full discrete Fourier transformation (DFT) (blue solid line), singular value decomposition (SVD) (magenta dashed line), and the nonlinear least square fit (LSF) (yellow dotted line). Results obtained with different sampling times are shown. The DFT needs a sampling time (395 fs) about 26 times larger than the SVD and LSF (15 fs).
@@ 3.2.1 Further numerical considerations
For the LSF to be accurate, one must carefully sample the key features of the signal, e.g. minima, maxima and turning points. This implies that, in general, the LSF needs more sampling points $P^{\alpha}_{i}$ (that is a higher sampling rate) than the DFT. This does not impact the cost of the simulations since a small time step is needed to integrate the equation of motion in Eq. (1). We found that non-uniform sampling (logarithmic or randomized) is more effective—thus reducing the sampling rate—than uniform sampling in capturing the key features of the signal, especially for high-frequency signals. An example of logarithmic sampling for LSF is shown in Fig. 3.
@@
Furthermore, a worst case for the LSF is when $\omega_{1}\approxeq\omega_{2}$. When they are exactly equal, the least square problem is ill-posed. For example, if we write down the first term of the second-order polarization:
$P^{(2)}$ $=\chi^{(2)}(2\omega_{1};\omega_{1},\omega_{1})\mathcal{E}^{2}(\omega_{1})+\chi^{(2)}(2\omega_{2};\omega_{2},\omega_{2})\mathcal{E}^{2}(\omega_{2})$
@@
and $\omega_{1}\approxeq\omega_{2}$ SHG and SFG become equivalent, their coefficients cannot be fixed using least square optimization. Similar issues arise if one of the frequency is an integer multiple of the other. In those cases, a few strategies may be adopted. One can eliminate the repeated terms in the fitting function, or use as starting values for the optimization those of the closest frequency already calculated, or simply interpolate the result from the neighbouring frequencies.
Note that the quasi-degenerate case is also the worst case for the DFT since closely spaced frequencies result in rapid beats in the signal. To accurately capture these beats, a very long sampling range is necessary to resolve the beat frequency ($\omega_{1}-\omega_{2}$). Ultimately, cases involving degenerate or quasi-degenerate frequencies are of little theoretical and experimental interest. Theoretically, for degenerate frequencies, the procedures in Ref. [11] (see Sec. 3.1) can extract the second-harmonic, while experimentally the interest lies in distinct frequencies resonant with different electron-hole excitations [8].
@@ 4 Computational Details
Ground-state properties of the h-BN mono- and bilayer and of MoS2 monolayer are calculated within the density functional theory using the Quantum-Espresso code[26]. We employ the Perdew-Burke-Ernzerhof (PBE) functional[27] with scalar-relativistic optimized norm-conserving pseudopotentials from the PseudoDojo repository (v0.4)[28] for the h-BN and FHI[29] for MoS2. The Kohn-Sham Hamiltonian is diagonalized for a given number of states (see $\bf{k}$-points and bands $N_{\mathrm{b}}$ in Table 1). These states are used as basis set to represent all operators that enter in the Eq. (1). All real-time simulations are carried out using the Yambo code [30]. The EOMs [Eq. (1)] are propagated using the Crank-Nicolson integrator with a time-step of 0.01 fs. In order to take dephasing into account, sampling was taken between 50-200 fs in all simulations. The static dielectric function $\epsilon_{\bf{G},\bf{G}^{\prime}}$ that enters in the calculation of screened-exchange self-energy, $\Sigma_{\text{SEX}}$ in Eq. (2), is calculated within the random-phase approximation (see Ref. [30] for more details). All calculations are performed in a supercell, so for each system, the susceptibility extracted from the time-dependent polarization is rescaled to the effective thickness, $d_{\text{eff}}$, resulting in $\chi_{\text{rescaled}}(\omega)=L_{z}/d_{\text{eff}}\cdot\chi(\omega)$ where $L_{z}$ labels the $z$ dimension of the supercell. To get the SFG and DFG spectra, we carry out simulations for all frequency couples, $\omega_{i},\omega_{j}$ in Eq. 3.2, in the desired energy ranges. We use the YamboPy code [31] to extract the relevant susceptibilities from the resulting polarizations [Eq. (4)], as detailed in Sec. 3. All parameters that enter in the different parts of the simulations are reported in Table 1.
@@ 5 Results
We apply the approach outlined in Secs. 2-3 to the SFG and DFG in h-BN and MoS2 monolayers (Secs. 5.1-5.2). h-BN monolayer is a wide band gap insulator with strong excitonic features and provides a clear example of the need for an accurate inclusion of excitonic effects. MoS2 monolayer is one of the most widely studied 2D material, including its SFG/DFG [4, 7]. For SFG and DFG, we show results in a heatmap, in which each point has been obtained by a separate real-time simulation. As a guide to reading such heatmaps (see e.g. Fig. 4), we can refer to Eqs. (5)-(6). The susceptibilities corresponding to SFG/DFG have poles both at one-photon ($\omega_{1,2}=\Omega_{\lambda}$) and at two-photon ($\omega_{1}\pm\omega_{2}=\Omega_{\lambda}$) resonances with single-particle transitions (or with excitons) in the system. The one-photon resonances correspond to vertical ($\omega_{1}$) and horizontal ($\omega_{2}$) lines in the SFG and DFG heatmaps. In the SFG heatmap, the two-photon resonances correspond to negative slope lines running from $\omega_{2}=\Omega_{\lambda}$ to $\omega_{1}=\Omega_{\lambda}$, while in DFG they correspond to positive slope lines starting at $\omega_{2}=\Omega_{\lambda}$ and $\omega_{1}=\Omega_{\lambda}$. Further, the diagonal $\omega_{1}=\omega_{2}$ corresponds to the SHG in the SFG and the optical rectification in the DFG heatmap. In Sec. 5.3, we consider THz-induced second-harmonic generation in 2L-h-BN. This system has inversion symmetry and thus has zero SHG at zero-field.

### v2

@@ 2.3 Field-induced second-harmonic generation
$\mathcal{P}(2\omega^{\pm})=\hat{\chi}^{(3)}(2\omega\pm\nu;\nu,\omega_{,}\omega)\mathcal{E}_{1}(\nu)\mathcal{E}^{2}_{2}(\omega).$ (9)
If $\mathcal{E}_{1}$ is static, i.e. $\nu=0$, the third-order susceptibility extracted from this contribution to the polarization gives the FI-SHG. Similarly, if $\nu\ll\omega$, i.e $\mathcal{E}_{1}$ is in the THz range, we have $2\omega\pm\nu\approx 2\omega$ and the expression in Eq. (9) can be rewritten as,
$\mathcal{P}(2\omega^{\pm})\approx\hat{\chi}^{(3)}(2\omega;\nu,\omega,\omega)\mathcal{E}_{1}(\nu)\mathcal{E}^{2}_{2}(\omega).$ (10)
Then, extracting the $\chi^{(3)}$ from $\mathcal{P}(2\omega^{+})$ or $\mathcal{P}(2\omega^{-})$, one obtains the corresponding FI-SHG for low-frequency time-dependent pump fields.
@@ 3 Signal processing: nonlinear susceptibilities
Adding an extra field to the formalism presented in Ref. [12] to obtain Eq. (1) is straightforward. The challenging part, and the main contribution of this work, is finding feasible, accurate strategies for extracting the relevant nonlinear susceptibilities from the resulting polarization $\bm{\mathcal{P}}(t)$. A strategy based on the discrete Fourier transform (FT) is used in Ref. [12] (Sec. 3.1) for the case of one external monochromatic external field. This analysis can be extended to more external monochromatic fields (Sec. 3.2), but as the common period for two, or more, (commensurate) frequencies can be of several hundreds of fs, this implies very long and computationally costly simulations. We thus propose an alternative strategy based on the least square fit (LSF), which turns out to be as accurate as the discrete FT without requiring too long simulations. The method proposed in this section is quite general and could be applied to Hamiltonians that differ from the one used in this work. Examples include tight-binding and other lattice models.
@@ 3.1 One external monochromatic field
$\bm{\mathcal{P}}(t)=\sum_{n=-\infty}^{\infty}\textbf{C}^{(n)}\exp[-i\omega^{(n)}t]\mathrm{,}$ (11)
with $\omega^{(n)}=n\omega_{L}$. The complex Fourier components $\textbf{C}^{(n)}$ can be determined by truncating the Fourier series to an order $S$ and sampling $2S+1$ values $\bm{\mathcal{P}}_{i}\equiv\bm{\mathcal{P}}(t_{i})$ within a period $T_{L}=2\pi/\omega_{L}$. For each direction $\alpha$, this yields
$\sum_{n=-S}^{S}\mathcal{F}^{(n)}_{i}C^{(n)}_{\alpha}=\mathcal{P}_{\alpha,i}\quad i=1,2S+1\mathrm{,}$ (12)
@@
Figure 2: SFG of h-BN with a pump frequency $\omega_{2}=3~\mathrm{eV}$ obtained at the IPA level using the full discrete Fourier transformation (FT) (blue solid line), singular value decomposition (SVD) (magenta dashed line), and the least square fit (LSF) (yellow dotted line). Results obtained with different sampling times are shown. The discrete FT needs a sampling time (385 fs) about 26 times larger than the SVD and LSF (15 fs). Below is displayed the difference in logarithmic scales for SVD and LSF, respectively, with 5 fs less sampling time to show the convergence.
@@ 3.2 Two external monochromatic fields
The resulting polarization $\bm{\mathcal{P}}(t)$ is periodic with period $T=2\pi/\omega_{0}$ and hence can be written as a Fourier series as in Eq. (11), in terms of the harmonics of $\omega_{0}$. As the frequencies of the external fields are multiple of the fundamental frequency $\omega_{1}=M\omega_{0}$ and $\omega_{2}=N\omega_{0}$, the $|M\pm N|$ harmonics correspond to the SFG/DFG. When it comes to setting up the system of linear equations (Eq. (12)), the sum over the harmonics must be truncated to an appropriate $S$ to include these processes (that is $S\gtrapprox M+N$). When compared with a single external field, the dimension of the system of linear equations is significantly larger. More critically, $T$ can be orders of magnitude larger than typical laser periods in the near-infrared to near-UV range ($\approx 1-5$ fs) for any pair of frequencies with more than one decimal leading to computationally intensive simulations (see Fig. 2).
Alternatively, we can expand the polarization $\bm{\mathcal{P}}(t)$ as the product of two Fourier series, one in terms of the harmonics of $\omega_{1}$ and the other in terms of the harmonics of $\omega_{2}$,
$\bm{\mathcal{P}}(t)=\sum_{n,m=-\infty}^{\infty}\bm{\mathrm{C}}^{(n,m)}\mathrm{exp}\left[-i(n\omega_{1}+m\omega_{2})t\right]\mathrm{,}$ (14)
@@
$\sum_{n=-S_{1}}^{S_{1}}\sum_{m=-S_{2}}^{S_{2}}\mathcal{F}_{i}^{(n,m)}C^{(n,m)}_{\alpha}=\mathcal{P}_{\alpha,i}\mathrm{,}$ (15)
where $\mathcal{F}_{i}^{(n,m)}\equiv\exp[-i(n\omega_{1}+m\omega_{2})t_{i}]$ and $S_{1},S_{2}$ are the maximum number of harmonics considered for each external field. Compared to the dicrete FT with one external field, it is straightforward to identify the relevant coefficients, for example, $n=1,m=1$ gives the SFG and $n=2,m=\pm 1$ the FI-SHG (choosing $\omega_{2}=\Omega$). On the other hand, generally, the system of linear equations in Eq. (15) is ill-conditioned. To solve Eq. (15), we calculate the Moore-Penrose inverse [26, 27], also called pseudoinverse, of $\mathcal{F}_{i}^{(n,m)}$ by using the singular value decomposition (SVD). This approach (that we will refer to as SVD) allows to work with a much smaller sampling time than $T$, so to significantly reduce the time from the simulations when compared with the discrete FT.
A further approach to obtain the Fourier coefficients is by solving a least squares problem [28]. This consists in finding the set of $C^{(n,m)}_{\alpha}$ that minimizes:
$\sum_{i=1}^{N}|\mathcal{P}_{\alpha}(t_{i})-\bar{\mathcal{P}}_{\alpha}(t_{i})|^{2},$ (16)
where $N$ is the number of sampling points and $\bar{\bm{\mathcal{P}}}$ is the Fourier series of Eq. (14) truncated to an order of $S$. Similarly to the SVD-based approach, the main advantage of the LSF over the discrete FT is that it is sufficient to sample part of the period $T$ to find the Fourier coefficient accurately. For instance, for h-BN, Fig. 3 compares the polarization from the simulation with that reconstructed from Eq. (14) with the coefficients obtained from LSF. Although only 15 fs are sampled, the LSF provides an accurate result.
The accuracy and performance of the three approaches are contrasted in Fig. 2 for the SFG of h-BN. As can be seen, SVD and LSF provide the same solution for the same sampling time due to the linear dependence between the components of polarization and Fourier matrix. In a more general nonlinear problem this is not necessarily the case. According to our tests, SVD and LSF are hence equally efficient. Using the approaches based on the SVD and the LSF allows to cut simulation time by a factor 26 compared to the discrete FT.
Figure 3: The time-dependent polarization (purple solid line) of h-BN calculated at the independent particle level with two electric fields ($\omega_{1}=0.2~\mathrm{eV}$, $\omega_{2}=3~\mathrm{eV}$). The signal can be divided into two regions: an initial “equilibration” region (up to $t\gg\gamma_{\mathrm{deph}}$, here $t=50~\mathrm{fs}$) during which the system’s eigenfrequencies are suppressed by dephasing and a region where Eq. (14) holds. In the second region the polarization is logarithmically sampled (teal dots) within a converged time window of 15 fs, smaller than the fundamental period of 20.678 fs of the signal. This sampling time is sufficient to correctly determine the Fourier coefficients by the least square fit (LSF) as verified by reconstructing the polarization (lavender dashed line) within the fundamental period using the Fourier coefficients and the truncated Eq. (14).
@@ 3.2.1 Further numerical considerations
For the LSF to be accurate, one must carefully sample the key features of the signal, e.g. minima, maxima, and turning points. This implies that, in general, the LSF needs more sampling points $\mathcal{P}^{\alpha}_{i}$ (that is a higher sampling rate) than the discrete FT. This does not impact the cost of the simulations since a small time step is needed to integrate the equation of motion in Eq. (1). We found that non-uniform sampling (logarithmic or randomized) is more effective—thus reducing the sampling rate—than uniform sampling in capturing the key features of the signal, especially for high-frequency signals. An example of logarithmic sampling for LSF is shown in Fig. 3.
@@
Furthermore, a worst case is when $\omega_{1}\approxeq\omega_{2}$. This increases the necessary simulation time a lot for the discrete FT. However, also for SVD and LSF this case is challenging as indicated in Fig. 2 by the peak of the logarithmic difference between different sampling times at around 3 eV ($\omega_{1}\approxeq\omega_{2}$). When they are exactly equal, the least square problem is ill-posed. For example, if we write down the first term of the second-order polarization:
$\mathcal{P}^{(2)}$ $=\chi^{(2)}(2\omega_{1};\omega_{1},\omega_{1})\mathcal{E}^{2}(\omega_{1})+\chi^{(2)}(2\omega_{2};\omega_{2},\omega_{2})\mathcal{E}^{2}(\omega_{2})$
@@
and set $\omega_{1}\approxeq\omega_{2}$, SHG and SFG become equivalent, i.e. their coefficients cannot be fixed using least square optimization. Similar issues arise if one of the frequencies is an integer multiple of the other. In those cases, a few strategies may be adopted. One can eliminate the repeated terms in the fitting function, or use as starting values for the optimization those of the closest frequency already calculated, or simply interpolate the result from the neighbouring frequencies.
Note that the quasi-degenerate case is also the worst case for the discrete FT since closely spaced frequencies result in rapid beats in the signal. To accurately capture these beats, a very long sampling range is necessary to resolve the beat frequency ($\omega_{1}-\omega_{2}$). Ultimately, cases involving degenerate or quasi-degenerate frequencies are of little theoretical and experimental interest. Theoretically, for degenerate frequencies, the procedures in Ref. [12] (see Sec. 3.1) can extract the second-harmonic, while experimentally the interest lies in distinct frequencies resonant with different electron-hole excitations [9].
@@ 4 Computational details
Ground-state properties of the h-BN mono- and bilayer and of MoS2 monolayer are calculated within the Density Functional Theory (DFT) using the Quantum-Espresso code [29]. We employ the Perdew-Burke-Ernzerhof (PBE) functional  [30] with scalar-relativistic optimized norm-conserving pseudopotentials from the PseudoDojo repository (v0.4) [31] for the h-BN and from Fritz-Haber Institute [32] for MoS2. The Kohn-Sham Hamiltonian is diagonalized for a given number of states (see $\bf{k}$-points and bands $N_{\mathrm{b}}$ in Table 1). These states are used as basis set to represent all operators that enter in Eq. (1). All real-time simulations are carried out using the Yambo code [21]. The EOMs [Eq. (1)] are propagated using the Crank-Nicolson integrator with a time-step of 0.01 fs. In order to take dephasing into account, sampling was taken between 50-200 fs in all simulations. The static dielectric function $\epsilon_{\bf{G},\bf{G}^{\prime}}$ that enters in the calculation of screened-exchange self-energy, $\Sigma_{\text{SEX}}$ in Eq. (4), is calculated within the random-phase approximation (see Ref. [21] for more details). All calculations are performed in a supercell, so for each system, the susceptibility extracted from the time-dependent polarization is rescaled to the effective thickness, $d_{\text{eff}}$, resulting in $\chi_{\text{rescaled}}(\omega)=L_{z}/d_{\text{eff}}\cdot\chi(\omega)$ where $L_{z}$ labels the $z$ dimension of the supercell. To get the SFG and DFG spectra, we carry out simulations for all frequency couples, $\omega_{i},\omega_{j}$, in the desired energy ranges. We use the YamboPy code [33] to extract the relevant susceptibilities from the resulting polarizations [Eq. (6)], as detailed in Sec. 5. All parameters that enter in the different parts of the simulations are reported in Table 1.
@@ 5 Results
We apply the approach outlined in Secs. 2-3 to the SFG and DFG in h-BN and MoS2 monolayers (Secs. 5.1-5.2). h-BN monolayer is a wide band gap insulator with strong excitonic features and provides a clear example of the need for an accurate inclusion of excitonic effects. MoS2 monolayer is one of the most widely studied 2D material, including its SFG/DFG [4, 7]. Since the h-BN and MoS2 monolayers belong to the $D_{3h}$ point group [34], they have only one non-vanishing second-order susceptibility tensor element with $\chi^{(2)}_{222}=-\chi^{(2)}_{211}=-\chi^{(2)}_{112}=-\chi^{(2)}_{121}\equiv\chi^{(2)}$ [^3] For MoS2 we have chosen the orientation of the $x$ axis aligned along an armchair direction following the conventions of Ref. [35].. For SFG and DFG, we show results in a heatmap, in which each point has been obtained by a separate real-time simulation. As a guide to reading such heatmaps (see e.g. Fig. 4), we can refer to Eqs. (7)-(8). The susceptibilities corresponding to SFG/DFG have poles both at one-photon ($\omega_{1,2}=\Omega_{\lambda}$) and at two-photon ($\omega_{1}\pm\omega_{2}=\Omega_{\lambda}$) resonances with single-particle transitions (or with excitons) in the system. The one-photon resonances correspond to vertical ($\omega_{1}$) and horizontal ($\omega_{2}$) lines in the SFG and DFG heatmaps. In the SFG heatmap, the two-photon resonances correspond to negative slope lines running from $\omega_{2}=\Omega_{\lambda}$ to $\omega_{1}=\Omega_{\lambda}$, while in DFG they correspond to positive slope lines starting at $\omega_{2}=\Omega_{\lambda}$ and $\omega_{1}=\Omega_{\lambda}$. Further, the diagonal $\omega_{1}=\omega_{2}$ corresponds to the SHG in the SFG and the optical rectification in the DFG heatmap. In Sec. 5.3, we consider THz-induced second-harmonic generation in 2L-h-BN. This system has inversion symmetry and thus has zero SHG at zero-field.